# PICT – A Differentiable, GPU-Accelerated Multi-Block PISO Solver for Simulation-Coupled Learning Tasks in Fluid Dynamics

### v1

@@ 1 Introduction
In this work, we introduce a novel implementation of a solver for incompressible fluid flows based on the PISO scheme called PICT, whose distinguishing feature is its differentiability. This allows gradients to flow through the solver, enabling end-to-end treatment of various optimization and machine learning (ML) tasks. Differentiable simulations have become increasingly prevalent in the field of robotics [9, 10, 11], and are garnering growing interest within the fluid dynamics community [12, 13, 14]. Differentiable solvers can be employed to optimize time-dependent problems step-by-step, a scenario that would be computationally very expensive to address with traditional adjoint formulations [15]. The use of differentiable solvers [9, 16, 17] in combination with machine learning algorithms can provide significant advantages. Typical supervised machine learning methods attempt to learn from examples and neural network models are optimized based on the gradient with respect to the mismatch between the network output and the reference ground truth. While these methods work well in tasks like replacing a numerical solver [18, 19] , super-resolution [20, 21] or flow field reconstruction [22, 23], trying to predict the temporal evolution of a physical system can lead to the accumulation of errors [24]. This issue can be addressed by integrating this evolution into the training via unrolling [25, 26], and by learning to predict consistent sequences. This concept has been successfully applied to turbulence modeling [27, 28, 29], for solver acceleration [30], and weather forecasting [31]. Similar to inductive biases like curl formulations [32], differentiable solvers inherently incorporate physical constraints into the optimization process, reducing the risk of physically implausible solutions. E.g. they can restrict correction terms to be divergence free in incompressible settings [27], wheres physics-informed losses would yield soft constrains [33]. In order to facilitate the integration of machine learning models in the simulations, we implement our solver in the widely used PyTorch framework [34]. It is available as open source software [^1] https://github.com/tum-pbs/PICT, leverages GPU acceleration, and provides custom gradient computations tailored to the solver.
@@
One application in which temporal consistency is particularly important is the development of learned turbulence closure models [35]. Different supervised approaches have been tested, both in the context of subgrid-scale (SGS) modeling [36, 37, 38] and wall-modeling. The latter approach has been implemented by training a classifier to choose among different models [39], or by using physics-informed neural networks [40]. Turbulence models for steady-state solutions via Reynolds-averaging have likewise been targeted with a PINN approach [41]. In order to address the problem of accumulating errors for transient problems, reinforcement learning (RL) based solutions to the turbulence modeling problem have been proposed [42, 43, 44, 45]. By formulating the problems as a Markov decision process (MDP), these approaches prevent the limitations of auto-regressive predictions in a purely supervised a-priori context. It should be noted, however, that the training procedure in this case requires extensive exploration of the solution space and induces high sample complexity. Differentiable solvers, on the other hand, directly compute gradients of the loss function with respect to system parameters, effectively yielding an a-posteriori training [46] and bypassing the need for iterative policy updates or extensive simulation runs. In this paper, we demonstrate the potential of our solver by training neural network models on wake and obstacle flows in 2D, and as subgrid-scale models for a turbulent channel flow (TCF) in 3D. Our paper is aligned with recent works that have investigated the use of differentiable solvers for SGS models using graph neural networks [47] and for shell models of turbulence [48]. A distinctive feature of our work is that the PICT solver allows for training the SGS model while only supervising in terms of velocity moments. I.e., no pre-computed training data sets are required in this case, and training only induces a moderate computational cost. The learned model significantly outperforms reference solvers in accuracy and runtime, and retains the target statistics for arbitrary timespans.
@@ 2.2 Multi-Block Grids and Transformations
We split the domain of interest into multiple blocks, based on the geometry of the boundaries, in order to obtain a multi-block grid, examples of which are shown in figure 2. Each block comprises a regular grid of quadrilateral (2D) or hexahedral (3D) elements that can be refined and aligned to boundaries with precomputed transformations. These transformations are represented as matrix $\mathbf{T}$ where the elements relate the Cartesian computational grid space with directions $\xi_{j}$ to the physical space with directions $x_{i}$ via $\mathbf{T}_{ji}=\partial\xi_{j}/\partial x_{i}$. Each side of a block can have a single boundary specified, either a connection to another block with matching resolution to create a conformal mesh or a prescribed quantity. Each block has separate velocity and pressure tensors that together make up the global fields. The linear systems for prediction and correction steps are assembled from the blocks and solved globally for the complete domain.
@@ 2.5 Operators for Turbulence Statistics
Production: $\begin{split}\text{Production: }P_{ij}&=-\left(\overline{u_{i}^{\prime}u_{k}^{\prime}}\frac{\partial\overline{u_{j}}}{\partial x_{k}}+\overline{u_{j}^{\prime}u_{k}^{\prime}}\frac{\partial\overline{u_{i}}}{\partial x_{k}}\right)\\ \text{Dissipation: }\epsilon_{ij}&=2\overline{\frac{\partial u_{i}^{\prime}}{\partial x_{k}}\frac{\partial u_{j}^{\prime}}{\partial x_{k}}}\\ \text{Turbulent transport: }T_{ij}&=-\frac{\partial\overline{u_{i}^{\prime}u_{j}^{\prime}u_{k}^{\prime}}}{\partial x_{k}}\\ \text{Viscous diffusion: }D_{ij}&=\frac{\partial^{2}\overline{u_{i}^{\prime}u_{j}^{\prime}}}{\partial x_{k}^{2}}\\ \text{Velocity pressure-gradient term: }\Pi_{ij}&=-\left(\overline{u_{i}^{\prime}\frac{\partial p}{\partial x_{j}}+u_{j}^{\prime}\frac{\partial p}{\partial x_{i}}}\right),\end{split}$ (9)
@@ 3 Optimization and Learning via Automatic Differentiation
The flexibility of our solver allows for the formulation of different initial and boundary condition optimization problems, as well as control tasks. In the present study, we focus our attention on developing correction models [25, 30] for highly under-resolved simulations. This means a neural network modifies the state of a simulation in order to conform to a learning task. In particular, our objective is to optimize deep neural networks $G(\cdot\,;\theta)$ parameterized by $\theta$ to output corrections that bring the simulation state closer to that of a higher-fidelity simulation. The output can be either a residual correction to the instantaneous velocity, $\mathrm{\mathbf{u}}_{\theta}:=G(\cdot\,;\theta)$ added to $\mathrm{\mathbf{u}}^{n}$ between simulation steps, or a correcting force added as additional source term $S_{\theta}:=G(\cdot\,;\theta)$ in eq. (1). While both are viable options, we focus on the latter in the following as it more tightly couples with the PISO solver.
@@ 3.1 Loss Terms and Physical Constraints
Alternatively, we may seek to minimize $\tilde{\mathrm{\mathbf{u}}}$ and thus to stabilize training by preventing the network from learning outputs that induce non-divergence-free motions [29]. A naive option consists in using the divergence of the velocity as a soft constraint of the form $||\nabla\cdot\mathrm{\mathbf{u}}_{\theta}||^{2}$. While this could drive the network towards divergence-free output, it is a local feedback for a global problem and thus could take many iterations to converge. Additionally, in our setting the central differencing used to compute the divergence would be prone to cause checkerboard artifacts [58]. A more principled alternative is to compute an additional pressure correction $\nabla^{2}p_{\theta}=\nabla\cdot\mathrm{\mathbf{u}}_{\theta}$. The spatial pressure gradient $\nabla p_{\theta}$ is then the exact, globally correct feedback that drives the velocity towards fulfilling the continuity equation, as $\mathrm{\mathbf{u}}_{\theta}-\nabla p_{\theta}$ represents a divergence free motion. To integrate this feedback we directly change the gradient $\partial L/\partial\theta$ to be computed as
@@ 3.2 Losses from Turbulence Statistics
$\begin{split}L_{U_{i}}^{n:m}&=\frac{1}{Y}\sum_{y=0}^{Y-1}||\overline{u_{i}}^{n:m}(y)-\overline{\hat{u}_{i}}(y)||_{2}^{2}\\ L_{u^{\prime}_{ij}}^{n:m}&=\frac{1}{Y}\sum_{y=0}^{Y-1}||\overline{u^{\prime}_{i}u^{\prime}_{j}}^{n:m}(y)-\overline{\hat{u}^{\prime}_{i}\hat{u}^{\prime}_{j}}(y)||_{2}^{2},\\ \end{split}$ (12)
where $Y$ is the resolution in wall normal direction and ¯$\overline{\Box}^{n:m}$ denotes averaging over homogeneous directions and rolled out steps $n$ to $m$. The loss terms for statistics are combined into a loss formulation that contains both temporally averaged statistics and per-frame statistics:
@@ 4 Validation
Figure 3: Velocity profiles for the 3D lid-driven cavity with $\text{Re}=1000$ for increasing resolutions. The left image of both pairs is the $u$-velocity on the vertical center line, and the right is the $v$-velocity on the horizontal center line. The left pair uses a uniform grid, the right a grid that was refined towards all boundaries. The velocities are normalized with the Reynolds number.
@@
Figure 5: A qualitative visualization of the pressure and velocity components of our TCF benchmark, a single wall is shown in computational space. Top: $u$ and $v$, bottom: $w$ and $p$.
In this section, we first give a brief overview of benchmark scenarios that were used to validate the forward simulations, focusing on 3D cases. Details and additional 2D validations are provided in B. We then validate the gradient derivations and investigate them in optimization settings.
@@ 4.1 Forward Simulation
For a lid-driven cavity setup, the plots in figure 3 show that the PICT solver correctly converges to the reference solution obtained from a high-resolution DNS [60]. Grid refinement, shown on the right of figure 3, further improves the results. We also simulate a 3D turbulent channel flow (TCF) at $\text{Re}_{\tau}=550$ and compare to established numerical references [59] and solvers [6, 61]. The resulting turbulence statistics, accumulated over 20 ETT after convergence of the simulation, can be found in figure 4. The flows are statistically stationary and the averaged, inner-scaled statistics are close to the spectral reference, despite the relatively low resolution. The resulting averaged $\text{Re}_{\tau}$ of 541 is also very close to the target. A qualitative visualization of the simulated boundary layer in terms of velocity components and pressure is shown in figure 5. The visualizations highlight the anisotropy of the velocity fields due to the presence of the wall and the highly directional nature of the flow, with the velocity streaks oriented in the streamwise direction.
@@ 4.2 Gradients
Figure 6: Loss curves for the optimization task and gradient path ablations (top row), and corresponding runtimes (bottom row). “Adv+P” denotes the full version, while “Adv” indicates that only the gradient of the advection-diffusion linear solve is used, “P” only that of the pressure linear solve. The “none” version relies on a gradient path without both linear solves. The columns show results with a rollout length of 1, 10, 100, and 100 (with lower learning rate) from left to right. The step size (equivalent to the learning rate) is set to $0.01$ for 1, 10, and 100 steps, and $0.001$ for the 100 steps on the right.
@@ 4.3 Gradient Path Ablation
Computationally, the most expensive parts of the simulation in terms are the linear solves for the advection-diffusion and pressure systems, which typically take 70-90% of the total runtime. While they are critical for the forward simulation, we investigated their influence on the gradient calculation, as alternative paths in the computational graph exist, cf. figure 1 and section 2.4. The versions we consider are Adv, where the gradients $\frac{\partial\mathrm{\mathbf{u}}^{*}}{\partial C}$ and $\frac{\partial\mathrm{\mathbf{u}}^{*}}{\partial{\mathrm{\mathbf{u}}}^{\text{RHS}}}$ of the advection solve are included, and P, utilizing $\frac{\partial p}{\partial P}$ and $\frac{\partial p}{\partial\nabla\cdot h}$. We then optimize the unknown scaling parameter for the initial velocity through a varying number of steps $n$ by computing the gradient $\partial\mathrm{\mathbf{u}}^{n}/\partial\mathrm{\mathbf{u}}^{0}$ via backpropagation through all $n$ steps of the solver. For shorter lengths of $n=1$ and 10, the difference is non-existent or negligible. However, when plotted against the wall clock time, there is a significant performance advantage in skipping the linear solves in the backwards pass, with a speed up of up to 2x when skipping both, as shown in figure 6. For larger $n$, on the other hand, the advection becomes more relevant, and when skipping both solves in the backwards pass the optimization can start to diverge. This happens in our setup when no linear solve gradients are used with $n=100$ steps and the default learning rate of $10^{-2}$. A significantly lower learning rate of $10^{-3}$, as used for the rightmost result shown in the figure, can compensate for this effect to some extend, but the optimization nonetheless converges to a sub-optimal minimum with a higher loss. In the 100 step case, the versions without the advection-diffusion solve gradients also no longer have a convergence-runtime advantage over the versions with these terms. Only the pressure solve gradient can be neglected, which still results in a faster convergence in the runtime comparisons, reaching the same loss in the same number of optimization steps.
@@
Table 1: Wall clock time to reach loss $<10^{-4}$ in seconds. All experiments run for 60 optimization steps with a learning rate of $lr=0.01$, except for $n=100$ which used 600 steps with $lr=0.001$.
@@ 5 Results - Deep Learning Applications
In this section we show the efficacy of our differentiable solver by developing learned correctors and SGS models to improve coarse simulations in various flow scenarios. As described in section 3, these models are neural networks $G(\cdot;\theta)$ that are tasked to estimate a correcting force $S_{\theta}$. We target two 2D scenarios, a vortex street and backward facing step, in addition to a turbulent channel flow in 3D. The cases highlight PICT’s support for non-uniform discretizations, as they employ grid refinement near obstacles and walls.
@@
The vortex shedding behavior transitions through distinct regimes depending on the Reynolds number. Laminar shedding typically occurs in the range $Re=40\sim 150$, transitioning to irregular, chaotic shedding between $Re=150\sim 300$. Beyond this, $Re>300$ enters the turbulent regime [69]. In this study, we focus on two cases with $Re=500$ and $Re=600$, for both training and testing. To introduce variability between training and test sets, the obstacle height $y_{s}$ and $Re$ are varied. A detailed breakdown of the geometry, Reynolds number, and sample ranges for each set is provided in table 2. For training and inference, we target a version down-sampled by 4$\times$ in the two spatial directions, giving a resolution of $67\times 36$, and a 10$\times$ larger timestep. We use an adaptive time stepping method to ensure CFL0.8absent0.8$\leq$ 0.8≤ 0.8 for all low- and high-resolution cases. Since the grids are not uniformly distributed with refinement, a coordinate-based approach that interpolates velocity values between the high- and low-resolution grids has been used to downsample high-resolution data. We employ a curriculum-based training strategy [25], starting with 4 unrolled steps and progressively increasing to 8 and finally 16 steps. The corrector models trained with 8 and 16 unrolled steps are referred to as $\text{NN}_{8}$ and $\text{NN}_{16}$, respectively. The training steps of $\text{NN}_{8}$ are extended to match $\text{NN}_{16}$ to ensure that the only variable differentiating $\text{NN}_{8}$ and $\text{NN}_{16}$ is the final number of unrolled steps. For training the NN corrector we use a simple MSE loss, which is evaluated at every other time step of the time integration. We simulate numerical solutions without a neural network at the same resolution, denoted as No-Model, which serves as a baseline for comparison against the trained models $\text{NN}_{8}$ and $\text{NN}_{16}$.
Dataset No. $y_{s}$ ($\mathrm{m}$) Re Sample range ($\mathrm{s}$) Unrolled steps
@@
Corr. MSE (×$\times 10^{-4}$) Corr. MSE (×$\times 10^{-3}$) Corr. MSE (×$\times 10^{-1}$)
@@
Since the stability of long-term forward simulations is inherently sensitive to initial conditions [70], we conducted a comprehensive assessment of the stability and accuracy improvements introduced by our methods. Initial states were sampled within a wide time range of $[70\,\mathrm{s},100\,\mathrm{s}]$, with an interval of $0.5\,\mathrm{s}$, resulting in a test set of 60 initial conditions. Each of them is advanced for $100\mathrm{s}$ in time, amounting to 2000 steps of the low-resolution simulator, which is far beyond the number of unrolled steps during training. In figure 7(a) we present a qualitative comparison of the time evolution of the different models. The baseline, No-Model, exhibits persistent checkerboard oscillations, particularly around the sharp corners of the obstacle. These artifacts remain present throughout the simulation, and the mismatch between the baselines and the reference solution grows over time. No-Model shows a fast growth of non-physical deterioration, with severe distortions in the vorticity field becoming evident by step 800. In contrast, both $\text{NN}_{8}$ and $\text{NN}_{16}$ suppress non-physical oscillations effectively, with no checkerboard patterns visible in the vorticity fields. The $\text{NN}_{16}$ model, by comparison, achieves superior performance. By step 1800, $\text{NN}_{16}$ produces a vorticity field nearly indistinguishable from the reference, showing its robustness and generalization capabilities for long-term simulations.
@@
This qualitative comparison provides an initial insight into the models’ effectiveness, which is further supported by quantitative evaluations, provided in figure 7(b) and table 3. The baseline shows a rapid decay in vorticity correlation with the reference solution within the first 100 steps. By step 2000, its correlation drops to a mean of 0.136, with significant variability across initial conditions. In contrast, both $\text{NN}_{8}$ and $\text{NN}_{16}$ perform significantly better in the initial 300 steps. However, beyond step 400, $\text{NN}_{8}$ experiences a decline in correlation. These deviations indicate that while $\text{NN}_{8}$ suppresses the numerical instability, it still suffers from a frequency decay during long-term forward simulations. This is caused by the limited number of unrolled steps in training, which are insufficient to fully capture the non-linear dynamics [26]. $\text{NN}_{16}$ demonstrates the best overall performance. It consistently maintains higher vorticity correlation throughout the simulation, achieving strong alignment with the reference solution, significantly outperforming the baselines and $\text{NN}_{8}$. This improvement highlights the importance and benefits of training via unrolling, as inherently supported by the PICT solver. It enables the model to better capture and mitigate long-term error accumulation.

### v2

@@ 1 Introduction
In this work, we introduce a novel implementation of a solver for incompressible fluid flows based on the PISO scheme called PICT, whose distinguishing feature is its differentiability. This allows gradients to flow through the solver, enabling end-to-end treatment of various optimization and machine learning (ML) tasks. Differentiable simulations have become increasingly prevalent in the field of robotics [9, 10, 11], and are garnering growing interest within the fluid dynamics community [12, 13, 14]. Differentiable solvers can be employed to optimize time-dependent problems step-by-step, a scenario that would be computationally very expensive to address with traditional adjoint formulations [15]. The use of differentiable solvers [9, 16, 17] in combination with machine learning algorithms can provide significant advantages. Typical supervised machine learning methods attempt to learn from examples and neural network models are optimized based on the gradient with respect to the mismatch between the network output and the reference ground truth. While these methods work well in tasks like replacing a numerical solver [18, 19], super-resolution [20, 21], flow field reconstruction [22, 23], or turbulent inflow generation [24], trying to predict the temporal evolution of a physical system can lead to the accumulation of errors [25]. This issue can be addressed by integrating this evolution into the training via unrolling [26, 27], and by learning to predict consistent sequences. This concept has been successfully applied to turbulence modeling [28, 29, 30], for solver acceleration [31], and weather forecasting [32]. Similar to inductive biases like curl formulations [33], differentiable solvers inherently incorporate physical constraints into the optimization process, reducing the risk of physically implausible solutions. E.g. they can restrict correction terms to be divergence free in incompressible settings [28], wheres physics-informed losses would yield soft constrains [34]. In order to facilitate the integration of machine learning models in the simulations, we implement our solver in the widely used PyTorch framework [35]. It will be made available as open source software [^1] https://github.com/tum-pbs/PICT, leverages GPU acceleration, and provides custom gradient computations tailored to the solver.
@@
One application in which temporal consistency is particularly important is the development of learned turbulence closure models [36]. Different supervised approaches have been tested, both in the context of subgrid-scale (SGS) modeling [37, 38, 39] and wall-modeling. The latter approach has been implemented by training a classifier to choose among different models [40], or by using physics-informed neural networks [41]. Turbulence models for steady-state solutions via Reynolds-averaging have likewise been targeted with a PINN approach [42]. In order to address the problem of accumulating errors for transient problems, reinforcement learning (RL) based solutions to the turbulence modeling problem have been proposed [43, 44, 45, 46]. By formulating the problems as a Markov decision process (MDP), these approaches prevent the limitations of auto-regressive predictions in a purely supervised a-priori context. It should be noted, however, that the training procedure in this case requires extensive exploration of the solution space and induces high sample complexity. Differentiable solvers, on the other hand, directly compute gradients of the loss function with respect to system parameters, effectively yielding an a-posteriori training [47] and bypassing the need for iterative policy updates or extensive simulation runs. In this paper, we demonstrate the potential of our solver by training neural network models on wake and obstacle flows in 2D, and as subgrid-scale models for a turbulent channel flow (TCF) in 3D. Our paper is aligned with recent works that have investigated the use of differentiable solvers for SGS models using graph neural networks [48] and for shell models of turbulence [49]. A distinctive feature of our work is that the PICT solver allows for training the SGS model while only supervising in terms of velocity moments. I.e., no pre-computed training data sets are required in this case, and training only induces a moderate computational cost. While the training process itself is stochastic, giving no formal guarantees on the exact results obtained, the learned model significantly outperforms reference solvers in accuracy and runtime, and retains the target statistics for arbitrary timespans and different initial conditions.
@@ 2.2 Multi-Block Grids and Transformations
As PICT is intended to be a general, differentiable solver for the incompressible Navier-Stokes equations, applicable to a broad range of learning tasks for fluid simulation, we choose a first order finite volume scheme for discretization. The domain of interest is further split into multiple blocks, based on the geometry of the boundaries, in order to obtain a multi-block grid, examples of which are shown in figure 2. Each block comprises a regular grid of quadrilateral (2D) or hexahedral (3D) elements that can be refined and aligned to boundaries with precomputed transformations. These transformations are represented as matrix $\mathbf{T}$ where the elements relate the Cartesian computational grid space with directions $\xi_{j}$ to the physical space with directions $x_{i}$ via $\mathbf{T}_{ji}=\partial\xi_{j}/\partial x_{i}$. Each side of a block can have a single boundary specified, either a connection to another block with matching resolution to create a conformal mesh or a prescribed quantity. Each block has separate velocity and pressure tensors that together make up the global fields. The linear systems for prediction and correction steps are assembled from the blocks and solved globally for the complete domain.
@@ 2.5 Operators for Turbulence Statistics
[…]
@@ 3 Optimization and Learning via Automatic Differentiation
The flexibility of our solver allows for the formulation of different initial and boundary condition optimization problems, as well as control tasks. In the present study, we focus our attention on developing correction models [26, 31] for highly under-resolved simulations. This means a neural network modifies the state of a simulation in order to conform to a learning task. In particular, our objective is to optimize deep neural networks $G(\cdot\,;\theta)$ parameterized by $\theta$ to output corrections that bring the simulation state closer to that of a higher-fidelity simulation. The output can be either a residual correction to the instantaneous velocity, $\mathrm{\mathbf{u}}_{\theta}:=G(\cdot\,;\theta)$ added to $\mathrm{\mathbf{u}}^{n}$ between simulation steps, or a correcting force added as additional source term $S_{\theta}:=G(\cdot\,;\theta)$ in eq. (1). While both are viable options, we focus on the latter in the following as it more naturally couples with the PISO solver.
@@ 3.1 Loss Terms and Physical Constraints
Alternatively, we may seek to minimize $\tilde{\mathrm{\mathbf{u}}}$ and thus to stabilize training by preventing the network from learning outputs that induce non-divergence-free motions [30]. A naive option consists in using the divergence of the velocity as a soft constraint of the form $\nabla\cdot\mathrm{\mathbf{u}}_{\theta}||^{2}$. While this could drive the network towards divergence-free output, it is a local feedback for a global problem and thus could take many iterations to converge. Additionally, in our setting the central differencing used to compute the divergence would be prone to cause checkerboard artifacts [59]. A more principled alternative is to compute an additional pressure correction $\nabla^{2}p_{\theta}=\nabla\cdot\mathrm{\mathbf{u}}_{\theta}$. The spatial pressure gradient $\nabla p_{\theta}$ is then the exact, globally correct feedback that drives the velocity towards fulfilling the continuity equation, as $\mathrm{\mathbf{u}}_{\theta}-\nabla p_{\theta}$ represents a divergence free motion. To integrate this feedback we directly change the gradient $\partial L/\partial\theta$ to be computed as
@@ 3.2 Losses from Turbulence Statistics
[…]
where $Y$ is the resolution in wall normal direction and $\overline{\Box}^{n:m}$ denotes averaging over homogeneous directions and rolled out steps $n$ to $m$. The loss terms for statistics are combined into a loss formulation that contains both temporally averaged statistics and per-frame statistics:
@@ 4 Validation
(a) Uniform grid
(b) Refined grid
Figure 3: Velocity profiles for the 3D lid-driven cavity with $\text{Re}=1000$ for increasing resolutions. The left image in each subfigure is the $u$-velocity on the vertical center line, and the right is the $v$-velocity on the horizontal center line. (b) uses a grid that was refined towards all boundaries. The velocities are normalized with the Reynolds number. The reference is a high-resolution DNS [60].
@@
Figure 5: Qualitative visualizations of the pressure and velocity components of our TCF benchmark. Only the flow close to the bottom wall in the computational space is shown. The color gradient and iso-surface values are chosen individually to highlight the detail of the flow structures. Top: $u$ and $v$, bottom: $w$ and $p$.
In this section, we first give a brief overview of benchmark scenarios that were used to validate the forward simulations, focusing on 3D cases. Details and additional validations are provided in B. We then validate the gradient derivations and investigate them in optimization settings. While PICT supports single and double precision, all experiments were run with single precision, matching the standard in learning applications, and conforming to recent analysis from turbulence modeling [62].
@@ 4.1 Forward Simulation
For a lid-driven cavity setup, the plots in figure 3 show that the PICT solver correctly converges to the reference solution obtained from a high-resolution DNS [60]. Grid refinement, shown on the right of figure 3, further improves the results. We also simulate a 3D turbulent channel flow (TCF) at $\text{Re}_{\tau}=550$ and compare to established numerical references [61] and solvers [6, 63]. The resulting turbulence statistics, accumulated over 20 ETT (eddy turnover time) after convergence of the simulation, can be found in figure 4. The flows are statistically stationary and the averaged, inner-scaled statistics of PICT closely match those obtained with OpenFOAM’s PISO implementation using the same mesh. Both solvers are also reasonably close to the spectral reference, despite the relatively low resolution. The resulting averaged $\text{Re}_{\tau}$ of 534 is also very close to the target. A qualitative visualization of the simulated boundary layer in terms of velocity components and pressure is shown in figure 5. The visualizations highlight the anisotropy of the velocity fields due to the presence of the wall and the highly directional nature of the flow, with the velocity streaks oriented in the streamwise direction.
@@ 4.2 Gradients
(a) $n=1$, lr =0.01=0.01
(b) $n=10$, lr =0.01=0.01
(c) $n=100$, lr =0.01=0.01
(d) $n=100$, lr =0.001=0.001
Figure 6: Loss curves for the optimization task and gradient path ablations (top row), and corresponding runtimes (bottom row). “Adv+P” denotes the full version, while “Adv” indicates that only the gradient of the advection-diffusion linear solve is used, “P” only that of the pressure linear solve. The “none” version relies on a gradient path without both linear solves. The columns show results with different rollout length $n$. The step size (equivalent to the learning rate) is additionally reduced from $0.01$ to $0.001$ for the longest rollout.
@@ 4.3 Gradient Path Ablation
Computationally, the most expensive parts of the simulation in terms are the linear solves for the advection-diffusion and pressure systems, which typically take 70-90% of the total runtime. While they are critical for the forward simulation, we investigated their influence on the gradient calculation, as alternative paths in the computational graph exist, cf. figure 1 and section 2.4. The versions we consider are Adv, where the gradients $\frac{\partial\mathrm{\mathbf{u}}^{*}}{\partial C}$ and $\frac{\partial\mathrm{\mathbf{u}}^{*}}{\partial{\mathrm{\mathbf{u}}}^{\text{RHS}}}$ of the advection solve are included, and P, utilizing $\frac{\partial p}{\partial P}$ and $\frac{\partial p}{\partial\nabla\cdot h}$. We then optimize the unknown scaling parameter for the initial velocity through a varying number of steps $n$ by computing the gradient $\partial\mathrm{\mathbf{u}}^{n}/\partial\mathrm{\mathbf{u}}^{0}$ via backpropagation through all $n$ steps of the solver. For shorter lengths of $n=1$ and 10, the difference is non-existent or negligible. However, when plotted against the wall clock time, there is a significant performance advantage in skipping the linear solves in the backwards pass, with a speed up of up to 2x when skipping both, as shown in figure 6. For larger $n$, on the other hand, the advection becomes more relevant, and when skipping both solves in the backwards pass the optimization can start to diverge. This happens in our setup when no linear solve gradients are used with $n=100$ steps and the default learning rate of $10^{-2}$. A significantly lower learning rate of $10^{-3}$, as used for the rightmost result shown in the figure, can compensate for this effect to some extent, but the optimization nonetheless converges to a sub-optimal minimum with a higher loss. In the 100 step case, the versions without the advection-diffusion solve gradients also no longer have a convergence-runtime advantage over the versions with these terms. Only the pressure solve gradient can be neglected, which still results in a faster convergence in the runtime comparisons, reaching the same loss in the same number of optimization steps.
@@
Table 1: Wall clock time to reach loss $10^{-4}$ in seconds. All experiments run for 60 optimization steps with a learning rate of $lr=0.01$, except for $n=100$ which used 600 steps with $lr=0.001$.
@@ 5 Results - Deep Learning Applications
In this section we show the efficacy of our differentiable solver by developing learned correctors and SGS models to improve coarse simulations in various flow scenarios. As described in section 3, these models are neural networks $G(\cdot;\theta)$ that are tasked to estimate a correcting force $S_{\theta}$. We target two 2D scenarios, a vortex street and backward facing step, in addition to a turbulent channel flow in 3D. The cases highlight PICT’s support for different learning formulations, matching both instantaneous flow fields and statistics, as well as non-uniform discretizations, as they employ grid refinement near obstacles and walls.
@@
[…]
Dataset No. $y_{s}$ Re Sample range Unrolled steps
@@
$C$ MSE ($\times 10^{-4}$) $C$ MSE ($\times 10^{-3}$) $C$ MSE ($\times 10^{-1}$)
@@
Since the stability of long-term forward simulations is inherently sensitive to initial conditions [75], we conducted a comprehensive assessment of the stability and accuracy improvements introduced by our methods. Initial states were sampled within a wide time range of $[70,100]$, with an interval of $0.5$, resulting in a test set of 60 initial conditions. Each of them is advanced by $100$ in time, amounting to 2000 steps of the low-resolution simulator, which is far beyond the number of unrolled steps during training. In figure 7(a) we present a qualitative comparison of the time evolution of the different models. The baseline, No-Model, exhibits persistent checkerboard oscillations, particularly around the sharp corners of the obstacle. These artifacts remain present throughout the simulation, and the mismatch between the baselines and the reference solution grows over time. No-Model shows a fast growth of non-physical deterioration, with severe distortions in the vorticity field becoming evident by step 800. In contrast, both $\text{NN}_{8}$ and $\text{NN}_{16}$ suppress non-physical oscillations effectively, with no checkerboard patterns visible in the vorticity fields. The $\text{NN}_{16}$ model, by comparison, achieves superior performance. By step 1800, $\text{NN}_{16}$ produces a vorticity field nearly indistinguishable from the reference, showing its robustness and generalization capabilities for long-term simulations.
This qualitative comparison provides an initial insight into the models’ effectiveness, which is further supported by quantitative evaluations, provided in figure 7(b) and table 3. Herein, the vorticity correlation is calculated as
$C(t)=\frac{\sum_{i,j}\omega_{i,j}(t)\cdot\hat{\omega}_{i,j}(t)}{\sqrt{\sum_{i,j}\left(\omega_{i,j}(t)\right)^{2}}\cdot\sqrt{\sum_{i,j}\left(\hat{\omega}_{i,j}(t)\right)^{2}}},$ (14)
@@
where $\omega_{i,j}(t)$ is the vorticity of the results and $\hat{\omega}_{i,j}(t)$ that of the reference at grid point $(i,j)$ and time $t$. The baseline shows a rapid decay in vorticity correlation with the reference solution within the first 100 steps. By step 2000, its correlation drops to a mean of 0.136, with significant variability across initial conditions. In contrast, both $\text{NN}_{8}$ and $\text{NN}_{16}$ perform significantly better in the initial 300 steps. However, beyond step 400, $\text{NN}_{8}$ experiences a decline in correlation. These deviations indicate that while $\text{NN}_{8}$ suppresses the numerical instability, it still suffers from a frequency decay during long-term forward simulations. This is caused by the limited number of unrolled steps in training, which are insufficient to fully capture the non-linear dynamics [27]. $\text{NN}_{16}$ demonstrates the best overall performance. It consistently maintains higher vorticity correlation throughout the simulation, achieving strong alignment with the reference solution, significantly outperforming the baselines and $\text{NN}_{8}$. This improvement highlights the importance and benefits of training via unrolling, as inherently supported by the PICT solver. It enables the model to better capture and mitigate long-term error accumulation. To further evaluate the performance of $\text{NN}_{16}$, we analyze the MSE of the aerodynamic coefficients (lift and drag), by comparing time-averaged values from the high-fidelity reference against both the no-model baseline and the $\text{NN}_{16}$ predictions. The baseline is numerically unstable, yielding a drag MSE of $2.14$. In contrast, our model ($\text{NN}_{16}$) successfully stabilizes the simulation, reducing the drag MSE by three orders of magnitude to $4.86\times 10^{-3}$. A clear improvement is also seen for the lift coefficient, where the MSE is reduced from $0.936$ in the unstable baseline to $1.29\times 10^{-3}$. This demonstrates the model’s success in capturing the fundamental flow dynamics where the baseline solver fails.